Round of the answer to this question to the nearest rand . David borrowed R911012 to refurbish his holiday home. The loan requires monthly repayment over 12 years . When he borrowed the money the interest rate was 12,4% per annum, compounded monthly but five years later the bank increase the annual interest rate to 13,9% in line with market rates . After five years the present value of the loan is R682081,77. With the new interest rate , his monthly payments will increase by ?


1
Expert's answer
2022-03-28T17:48:49-0400

If the present value of the loan is R911012 to be repaid after 12 years with an interest rate of 12.4% but to be changed to a new rate of 13.9% after the fifth year of starting, the monthly payments for the first five years is:

"Pmt=\frac{PV}{[\frac{1-(1+\frac{r}{m})^{-(mn)}}{\frac{r}{m}}]}"

"Pmt=\frac{R911012}{[\frac{1-(1+\frac{12.4\%}{12})^{-(12×5)}}{\frac{12.4\%}{12}}]}"

"Pmt=R20450"


After the fifth year, the interest rate changes to 13.9%, the present value becomes R682018.77 and n becomes 7

"\therefore Pmt=\frac{R682081.77}{[\frac{1-(1+\frac{13.9\%}{12})^{-(12×7)}}{\frac{13.9\%}{12}}]}"

"Pmt=R12745"


David's monthly payments decreases by R7705 (R20450-R12745).